1. Twenty-five student were asked the question How many times per week do you eat chips? The following results were found: # of Times 0 8 4 Construct a Probability Distribution for the random variable x, where x is the number of times per week a student eats chips. A) B) Find the probability someone eats chips at most 2 times per week. C) Find the probability someone eats chips more than 3 times per week. D) Find the probability someone eats chips at least 3 but no more than 4 times per week E) Find the mean number of times someone eats chips per week.
Answer #1

A)

X P(X)
0 7/25 =0.28
1 8/25 = 0.32
2 3/25 = 0.12
3 2/25 = 0.08
4 2/25 = 0.08
5 3/25 = 0.12

B) P(at most 2) = 0.28+0.32+0.12 = 0.72

C) P(more than 3) = 0.08 + 0.12 = 0.20

D) P(at least 3 but not more than 4) = 0.08+0.08 = 0.16

E) Mean = Sum of {XP(X)}

= 0x0.28 + 1x0.32 + 2x0.12 + 3x0.08 + 4x0.08 + 5x0.12

= 1.72